OpenStudy (anonymous):

please help! will give medal! which is equal to 3^3 x 3^5

OpenStudy (imstuck):

Is that \[3^{3}\times 3^{5}\]?

OpenStudy (anonymous):

yes

jhonyy9 (jhonyy9):

so us a^x *a^y = a^(x+y)

OpenStudy (anonymous):

270

OpenStudy (imstuck):

ok, just like with any other exponential expressions, as long as the base is the same (and it is...it's 3), you add the exponents and raise that base to that power.

OpenStudy (anonymous):

answer is 270

OpenStudy (anonymous):

can i get medal

OpenStudy (anonymous):

oh okay thank you

OpenStudy (anonymous):

can i get medal

OpenStudy (imstuck):

how about maybe @Darknight200 NOT just gives out free answers for medals. The point is to HELP

OpenStudy (imstuck):

And TY for the medals~!~ ; )

OpenStudy (anonymous):

true im srry

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):

you would do 3 times 3 times 3 and solve it

OpenStudy (anonymous):

then you would do 3 times 3 times 3 times 3 times 3 and solve it

OpenStudy (anonymous):

then add them

OpenStudy (anonymous):

its asking which is it equal to

OpenStudy (anonymous):

yes and it equals 270

OpenStudy (anonymous):

but the only answers it has is \[3^{8}\] and stuff like that

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

i think i know what the answer is though

OpenStudy (anonymous):

that is the answer 8 3

OpenStudy (anonymous):

ok wellc can i have a medal

OpenStudy (anonymous):

i gave it to imstuck because he explained it a little better you just gave the answer but thanks for your help

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):

even tho i dont have any medals

OpenStudy (anonymous):

wanna be friends

OpenStudy (anonymous):

fan me

OpenStudy (anonymous):

didi

OpenStudy (anonymous):

and ill fan you

OpenStudy (anonymous):

k i did but i have to go i will talk to you later

OpenStudy (anonymous):

k